# WorldPrompter: Traversable Text-to-Scene Generation

### v1

@@ Abstract
Scene-level 3D generation is a challenging research topic, with most existing methods generating only partial scenes and offering limited navigational freedom. We introduce WorldPrompter, a novel generative pipeline for synthesizing traversable 3D scenes from text prompts. We leverage panoramic videos as an intermediate representation to model the 360° details of a scene. WorldPrompter incorporates a conditional 360° panoramic video generator, capable of producing a 128-frame video that simulates a person walking through and capturing a virtual environment. The resulting video is then reconstructed as Gaussian splats by a fast feedforward 3D reconstructor, enabling a true walkable experience within the 3D scene. Experiments demonstrate that our panoramic video generation model achieves convincing view consistency across frames, enabling high-quality panoramic Gaussian splat reconstruction and facilitating traversal over an area of the scene. Qualitative and quantitative results also show it outperforms the state-of-the-art 360° video generators and 3D scene generation models.
@@
Figure 1: We present WorldPrompter, a method that generates traversable environments from a text prompt alone (top). Our method first generates a spatially coherent video (2nd row) that travels through the environment described by the prompt. Next, we elevate the scene into a 3D Gaussian Splat (bottom), from which we project some views into standard perspective images (3rd row).
*Work done as an intern at Adobe Research.
1 Introduction
@@
Scene-level 3D generation has attracted increasing research interest in recent years. For example, methods like RealmDreamer [38] and LucidDreamer [10] can generate Gaussian Splatting (GS) 3D scenes from text prompts. They rely on single-view diffusion models and monocular depth estimation to generate scenes progressively from one view to another by 2D inpainting. However, the 3D scenes created by these methods are confined to a limited viewable area and are not freely traversable. This constraint differentiates these approaches from a ”complete” scene generation that offers a navigable virtual world, similar to what is found in interactive video games.
@@
In this paper, we propose WorldPrompter, the first generative pipeline that produces a navigable 3D scene from text prompts. Training a model to directly synthesize a full scene can be very challenging due to highly scarce scene-level 3D data. Drawing inspiration from conventional 3D capture methods and recent advancements in video generation models ([32, 54, 4]), we propose using videos as an intermediate representation to model complete 3D environments. Specifically, we train a diffusion model to generate videos traversing the environment based on text prompts. Moreover, instead of synthesizing captures from perspective cameras, we leverage panoramic cameras that directly capture the entire surroundings of an environment, providing 360° viewpoints that are ideal for our application.
@@
Specifically, WorldPrompter is a two-stage pipeline composed of a 360° panoramic video generation model and a 3D reconstruction step to generate the final 3D Gaussians, a summary of which is shown in Fig. 1. The text-conditioned panoramic video generation model is trained to generate a 360° panoramic video that simulates a person walking through an environment while holding a panoramic camera to capture the scene. To train the video generator, we curated a dataset comprising approximately 1,700 panoramic videos and a panorama image dataset containing more than 200,000 high-quality panoramic in-the-wild captures. To enhance the generative capability of our method, we fine-tuned a Diffusion Transformer (DiT)-based video generator on our dataset to produce 10.6-second 360° panoramic videos (128 frames at 12 fps). We introduce a masked diffusion loss to ensure a clean panoramic capture sequence free from the person holding the camera, ready for subsequent reconstruction purposes. From the generated videos, we extract panoramic frames and project them from their original spherical coordinates onto the image plane to obtain perspective images from multiple viewpoints throughout the scene. The reconstruction step then builds the final 3D Gaussian Splat (3DGS) scene from these cropped images using a fast feedforward reconstruction model [56]. The resulting scene allows users to navigate freely within a virtual environment.
We show that our 360° video generation model achieves state-of-the-art generation capability in view coverage, temporal consistency, and visual quality. The high view consistency enables the videos to be reliably used as proxies for 3D scene generation, resulting in a significantly better view coverage of the generated scene and a traversable scene, more flexible than prior art. We highlight our contributions as follows:
  • We present WorldPrompter, the first generative pipeline to synthesize a complete 3D scene from a text prompt.
  • We propose leveraging panoramic capture videos as intermediate proxies for full 3D scene generation, addressing the scarcity of complete 3D scene datasets.
  • We introduce a high-quality 360° panoramic video generation model trained with mixed datasets and a masked diffusion loss achieving superior temporal consistency and view coverage compared to previous work.
  • We show that our 360° video is consistent enough to be efficiently turned into a 3DGS scene.
@@
Figure 2: Overview of WorldPrompter. (top) We train a text-to-video model on a mix of 360° videos and images depicting in-the-wild environments. As it is challenging to avoid the person and camera equipment being visible in the video capture, we mask these elements out of the frame using a pretrained image segmentation model [1], and obtain the prompts from the video frames using LLaVA [24] for the videos, and BLIP-2 [19] for images. (bottom) At inference time, a user supplies a text prompt to our text-to-video model, which produces a “walk-through” video of the scene, which we reconstruct into a 3D Gaussian splat representation using Long-LRM [57].
2 Related Work
2.1 Panoramic Image and Video Generation
@@
In computer graphics, 360° panoramas—often called IBLs or environment maps—are widely used to represent distant lighting and environments [12]. Numerous studies estimate lighting as High Dynamic Range (HDR) panoramas from input photos [51, 44, 11]. However, the resolution of these panoramas is often limited, making them suitable for light representation but insufficient for 3D reconstruction.
@@
Advancements in diffusion models have made high-resolution image synthesis possible [35, 13]. Panorama generation also benefits from the use of diffusion models. For example, Text2Light [9] generates 4K HDR panoramas via text-driven low-resolution generation and super-resolution inverse tone mapping. MVDiffusion [39] creates panoramas by stitching consistent multi-view images generated from text prompts using attention layers. PanFusion [52] employs a dual-branch diffusion model with equirectangular-Perspective Projection Attention (EPPA) to improve consistency and control across panorama and perspective domains.
@@
Diffusion models have also demonstrated their potential in video generation [6, 5]. Diffusion Transformers [31] offer a more versatile method for modeling high-dimensional data, making them particularly well-suited for video generation [7]. For 360° video generation, Wang et al. propose 360DVD for generating 360° panoramic videos from text prompts and motion conditions, where they leverage a lightweight 360-Adapter to finetune the Stable Diffusion [35] text-to-image model. However, their model is limited to generating only 16 frames, whereas our 360° video generation model produces significantly more frames with superior temporal consistency.
2.2 3D Neural Representation and Reconstruction
3D reconstruction and representation is an extensively studied field. While mesh-based methods [17] and point-based methods [33] have been extensively explored, they are limited in the quality of their results. In contrast, neural representations offer a flexible way to reconstruct 3D data. For instance, Neural Radiance Fields (NeRF) [26] introduces a novel way to represent 3D scenes by encoding volumetric scene information as a continuous function parameterized by a neural network, which inspired many follow-up works [27, 2, 3] that optimize efficiency, quality, and rendering speed. However, the MLP-based paradigm requires training a distinct model per scene, making reconstruction time-consuming. Additionally, rendering views involves model queries, creating a bottleneck for real-time applications.
@@
3D Gaussian Splatting (3DGS) [18] introduces a memory-efficient and low-overhead 3D implicit representation using Gaussians derived from Structure-from-Motion (SfM) points. By combining anisotropic covariance optimization with a fast visibility-aware rendering algorithm, 3DGS enables efficient scene representation, high-quality real-time rendering, and has quickly emerged as a new paradigm in neural reconstruction [49, 47, 16, 14, 8]. Nevertheless, 3DGS still requires an optimization process, which can be costly for larger scenes. Recent advances such as Large Reconstruction Models (LRMs) [56, 45, 48, 53, 41] target scene reconstruction with a single feed-forward evaluation during inference. They achieve a reconstruction quality comparable to optimization-based methods while significantly improving their efficiency. In our scene generation pipeline, we also utilize this state-of-the-art feed-forward reconstruction approach for long-sequence reconstruction, achieving fast reconstruction.
2.3 3D Scene Generation
@@
3D generation has emerged as a prominent focus in generative modeling. Given the scarcity of high-quality 3D data, recent works usually leverage intermediate representations, such as multi-view images, to improve the quality and efficiency of their reconstructions. Many studies have explored generating 3D objects in various formats, including Gaussian splats and meshes. For instance, Instant3D [20] accelerates 3D asset creation from text prompts by combining sparse-view generation with transformer-based reconstruction. Similarly, MVDream [37] improves consistency through multi-view diffusion by integrating 2D and 3D data. One-2-3-45 [25] simplifies single-image 3D reconstruction, enabling view-consistent mesh generation.
@@
Scene-level generation, however, is a more challenging task because it requires more observations to achieve comprehensive view coverage. Typically, 3D scene generation methods employ progressive pipelines [29, 10, 15]: anchor cameras are first selected to generate text-aligned 2D images and corresponding depth maps, followed by a neural reconstruction process to build the initial 3D structure. This process is iteratively repeated to expand scene coverage by generating additional camera views and inpainting missing regions. However, such methods often suffer from prolonged generation times and incomplete coverage. Other studies have explored interactive and layout-guided approaches for 3D scene generation. Systems like WonderWorld [50] provide user-friendly interfaces for scene customization, while Director3D [22] facilitates real-world scene creation with adaptive camera trajectories.
@@
The closest method to ours is DreamScene360 [55], which generates 3D scenes from text prompts using a single panorama image, along with depth estimation and Gaussian Splats optimization. However, while the resulting scene supports 360° viewing, it suffers from a limited movable range. 4K4DGen [21], introduces panorama-to-4D generation but still remains limited in terms of traversable area. More recently, ODIN [40] proposed learning 3D from 360° videos. However, their approach extracts only cropped views from videos to perform image-to-image style novel view synthesis, under-utilizing the full potential of 360° videos. As a result, ODIN fails to model a complete 3D world and can only generate partial scenes. In contrast to these approaches, our model generates complete scenes by fully leveraging the 360° representation, ensuring consistency and comprehensive coverage through a panoramic video generation model.
3 WorldPrompter
3.1 Overview
@@
Our WorldPrompter synthesizes a traversable, 360° viewable 3D scene in two stages: generation and reconstruction. The generation stage leverages a text-conditioned 360° panoramic video diffusion model to generate a 10.6-second panoramic capture video for a target scene, simulating someone holding the camera and walking through the environment. Later, the reconstruction stage runs COLMAP [36] to estimate precise camera poses from perspective crops of the generated videos and builds a navigable 3DGS using Long-LRM [56] with the calibrated camera poses. The generated panoramic video features rich details and geometrically consistent visual details from multiple viewpoints along the traversal trajectory, ensuring high-quality reconstruction by the reconstruction step and yielding appealing novel views in the traversable scene.
3.2 Video Generator for 360° Panoramic Captures

### v4

# Generating 360° Video is What You Need For a 3D Scene

@@ Abstract.
Generating 3D scenes is still a challenging task due to the lack of readily available scene data. Most existing methods only produce partial scenes and provide limited navigational freedom. We introduce a practical and scalable solution that uses 360° video as an intermediate scene representation, capturing the full-scene context and ensuring consistent visual content throughout the generation. We propose WorldPrompter, a generative pipeline that synthesizes traversable 3D scenes from text prompts. WorldPrompter incorporates a conditional 360° panoramic video generator, capable of producing a 128-frame video that simulates a person walking through and capturing a virtual environment. The resulting video is then reconstructed as Gaussian splats by a fast feedforward 3D reconstructor, enabling a true walkable experience within the 3D scene. Experiments demonstrate that our panoramic video generation model, trained with a mix of image and video data, achieves convincing spatial and temporal consistency for static scenes. This is validated by an average COLMAP matching rate of 94.6%, allowing for high-quality panoramic Gaussian splat reconstruction and improved navigation throughout the scene. Qualitative and quantitative results also show it outperforms the state-of-the-art 360° video generators and 3D scene generation models.
Diffusion Models, 360° Video Generative Model, 3D Reconstruction
submissionid: 1445journal: TOG
@@
Figure 1. We present WorldPrompter, a method that generates traversable environments from a text prompt alone (above). Our method first generates a spatially coherent 360° video capture (top) that travels through a static environment described by the prompt. Next, we elevate generated video capture into a 3D Gaussian Splat (middle), enabling free navigation within a 3D scene using a virtual perspective camera to generate novel views (bottom).
1. Introduction
@@
Scene-level 3D generation has attracted increasing research interest in recent years. The limited availability of data generally hinders direct 3D scene generation (Liu et al., 2024a), making it difficult for methods to generalize to realistic scene synthesis effectively. To address this challenge, recent approaches typically rely on an image diffusion model to create realistic scene views. For example, methods like RealmDreamer (Shriram et al., 2025) and LucidDreamer (Chung et al., 2023) can generate Gaussian Splatting (GS) 3D scenes from text prompts. They rely on single-view diffusion models and monocular depth estimation to generate scenes progressively from one view to another by 2D inpainting. However, the 3D scenes created by these methods are confined to a limited viewable area and are not freely traversable. This constraint differentiates these approaches from a ”complete” scene generation that offers a navigable virtual world, similar to what is found in interactive video games. Recent works (Zhou et al., 2024; Yang et al., 2024) attempt to bridge the gap between partial and full scene generation by leveraging panoramic images. However, these methods still rely on outpainting to fill in missing regions, which inherently limits the navigable range and often leads to distortions during viewpoint transitions.
More importantly, we observe that scenes generated via these outpainting-based approaches frequently suffer from distortion and inconsistencies in visual content across the scene. We hypothesize that these issues stem from the commonly used progressive outpainting approach, which independently synthesizes new content based on the previously generated content. However, this assumption is flawed, as visual content is globally interconnected and consistent rather than strictly sequential (Tian et al., 2024). As a result, while their generated scenes may appear locally plausible, they exhibit significant global distortions (Fig. 7).
@@
In this paper, instead of synthesizing a scene in a view-by-view fashion, we aim to generate a maximally navigable scene in a single pass, ensuring undistorted global visual content synthesis. Although scenes can be represented in various formats, such as meshes or point clouds, data in these formats is often limited and difficult to scale. Drawing inspiration from conventional 3D capture methods and recent advancements in video generation models ((Polyak et al., 2024; Zheng et al., 2024; Blattmann et al., 2023a)), we propose using videos as an intermediate representation to model complete 3D environments. Modeling scenes using video captures provides a scalable approach for creating training data, enabling generative models to achieve strong generalization capabilities. To validate this idea, we propose WorldPrompter, a generative pipeline that produces a navigable 3D scene from text prompts. Specifically, we train a diffusion model to generate videos traversing the environment based on text prompts. Moreover, instead of synthesizing captures from perspective cameras, we leverage panoramic cameras that directly capture the entire surroundings of an environment, providing 360° viewpoints that are ideal for our application.
@@
WorldPrompter is a two-stage pipeline composed of a 360° panoramic video generation model and a 3D reconstruction step to generate the final 3D Gaussians, a summary of which is shown in Fig. 1. The text-conditioned panoramic video generation model is trained to generate a 360° panoramic video that simulates a person walking through an environment while holding a panoramic camera to capture the scene. To train the video generator, we collected a dataset comprising approximately 1,700 panoramic videos and a panorama image dataset containing more than 200,000 high-quality panoramic in-the-wild captures. To enhance the generative capability of our method, we fine-tuned a Diffusion Transformer (DiT)-based video generator on our dataset to produce 10.6-second 360° panoramic videos (128 frames at 12 fps). We introduce a masked diffusion loss to ensure a clean panoramic capture sequence free from the person holding the camera, ready for subsequent reconstruction purposes. From the generated videos, we extract panoramic frames and project them from their original spherical coordinates onto the image plane to obtain perspective images from multiple viewpoints throughout the scene. The reconstruction step then builds the final 3D Gaussian Splat (3DGS) scene from these cropped images using a fast feedforward reconstruction model (Ziwen et al., 2024). The resulting scene allows users to navigate freely within a virtual environment, beyond the original video camera path.
We demonstrate that our 360° video generation model—trained on a mixture of image and video data—achieves state-of-the-art performance in temporal consistency and visual quality. The high consistency of the views, confirmed by a matching rate of up to 94.5% using COLMAP (Schönberger and Frahm, 2016), ensures that the videos can be reliably used as proxies for 3D scene generation. This results in a traversable, 360° viewable environment with globally coherent visual content. We demonstrate that our generated scene allows longer navigation range and offers superior global visual coherence and quality compared to prior art. We highlight our contributions as follows:
  • We propose leveraging panoramic capture videos as an intermediate 3D scene representation. We found it is a practical solution for scaling up training data for high-quality photorealistic scene synthesis.
  • We present WorldPrompter, a practical generative text-to-scene pipeline that synthesizes a traversable 3D scene with strong global visual coherence.
  • We introduce a high-quality 360° panoramic video generation model, where the proposed mixed datasets and masked diffusion loss play a critical role in achieving artifact-free and generalizable static 360° scene generation.
@@
Figure 2. Overview of WorldPrompter. (top) We train a text-to-video model on a mix of 360° videos and images depicting in-the-wild environments. As it is challenging to avoid the person and camera equipment being visible in the video capture, we mask these elements out of the frame using a pretrained image segmentation model (Adobe, 2025), and obtain the prompts from the video frames using LLaVA (Liu et al., 2023b) for the videos, and BLIP-2 (Li et al., 2023a) for images. (bottom) At inference time, a user supplies a text prompt to our text-to-video model, which produces a “walk-through” video of the scene, which we reconstruct into a 3D Gaussian splat representation using Long-LRM (Ziwen et al., 2024).
2. Related Work
2.1. Panoramic Image and Video Generation
@@
In computer graphics, 360° panoramas—often called IBLs or environment maps—are widely used to represent distant lighting and environments (Debevec and Malik, 2023). Numerous studies estimate lighting as High Dynamic Range (HDR) panoramas from input photos (Zhan et al., 2021; Weber et al., 2022; Dastjerdi et al., 2023). However, the resolution of these panoramas is often limited, making them suitable for light representation but insufficient for 3D reconstruction.
@@
Advancements in diffusion models have made high-resolution image synthesis possible (Rombach et al., 2022; Esser et al., 2024). Panorama generation also benefits from the use of diffusion models. For example, Text2Light (Chen et al., 2022) generates 4K HDR panoramas via text-driven low-resolution generation and super-resolution inverse tone mapping. MVDiffusion (Tang et al., 2023) creates panoramas by stitching consistent multi-view images generated from text prompts using attention layers. PanFusion (Zhang et al., 2024b) employs a dual-branch diffusion model with equirectangular-Perspective Projection Attention (EPPA) to improve consistency and control across panorama and perspective domains. DiffPano (Ye et al., 2024a) fine-tunes Stable Diffusion (Rombach et al., 2022) using LoRA (Hu et al., 2021) to generate panorama images, and with a Spherical Epipolar-Aware Attention Module, it enables multi-view panorama generation.
@@
Diffusion models have also demonstrated their potential in video generation (Blattmann et al., 2023b, a). Diffusion Transformers (Peebles and Xie, 2023) offer a more versatile method for modeling high-dimensional data, making them particularly well-suited for video generation (Brooks et al., 2024). For 360° video generation, Wang et al. propose 360DVD for generating 360° panoramic videos from text prompts and motion conditions, where they leverage a lightweight 360-Adapter to finetune the Stable Diffusion (Rombach et al., 2022) text-to-image model. However, their model is limited to generating only 16 frames, whereas our 360° video generation model produces significantly more frames with superior temporal consistency.
2.2. 3D Neural Representation and Reconstruction
@@
3D reconstruction and representation is an extensively studied field. While mesh-based methods (Kazhdan et al., 2006) and point-based methods (Qi et al., 2017) have been extensively explored, they are limited in the quality of their results. In contrast, neural representations offer a flexible way to reconstruct 3D data. For instance, Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) introduces a novel way to represent 3D scenes by encoding volumetric scene information as a continuous function parameterized by a neural network, which inspired many follow-up works (Müller et al., 2022; Barron et al., 2022, 2023) that optimize efficiency, quality, and rendering speed. However, the MLP-based paradigm requires training a distinct model per scene, making reconstruction time-consuming. Additionally, rendering views involves model queries, creating a bottleneck for real-time applications. 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) introduces a memory-efficient and low-overhead 3D implicit representation using Gaussians derived from Structure-from-Motion (SfM) points. By combining anisotropic covariance optimization with a fast visibility-aware rendering algorithm, 3DGS enables efficient scene representation, high-quality real-time rendering, and has quickly emerged as a new paradigm in neural reconstruction (Ye et al., 2024b; Wu et al., 2024; Jiang et al., 2024; Gao et al., 2024; Charatan et al., 2024). Nevertheless, 3DGS still requires an optimization process, which can be costly for larger scenes. Recent advances such as Large Reconstruction Models (LRMs) (Ziwen et al., 2024; Wei et al., 2024; Xie et al., 2024; Zhang et al., 2024a; Wang et al., 2023) target scene reconstruction with a single feed-forward evaluation during inference. They achieve a reconstruction quality comparable to optimization-based methods while significantly improving their efficiency. In our scene generation pipeline, we also utilize this state-of-the-art feed-forward reconstruction approach for long-sequence reconstruction, achieving fast reconstruction.
2.3. 3D Scene Generation
@@
3D generation has emerged as a prominent focus in generative modeling. Given the scarcity of high-quality 3D data, recent works usually leverage intermediate representations, such as multi-view images (Liu et al., 2023c; Voleti et al., 2024; Hu et al., 2024; Li et al., 2023b; Shi et al., 2024; Liu et al., 2024b) and videos (Parthasarathy et al., 2024; Han et al., 2024; Chen et al., 2024), to improve the quality and efficiency of their reconstructions. Many studies have explored generating 3D objects in various formats, including Gaussian splats and meshes. For instance, Instant3D (Li et al., 2023b) accelerates 3D asset creation from text prompts by combining sparse-view generation with transformer-based reconstruction. Similarly, MVDream (Shi et al., 2024) improves consistency through multi-view diffusion by integrating 2D and 3D data. One-2-3-45 (Liu et al., 2024b) simplifies single-image 3D reconstruction, enabling view-consistent mesh generation.
@@
Scene-level generation, however, is a more challenging task because it requires more observations to achieve comprehensive view coverage. Typically, 3D scene generation methods employ progressive pipelines (Ouyang et al., 2023; Chung et al., 2023; Höllein et al., 2023): anchor cameras are first selected to generate text-aligned 2D images and corresponding depth maps, followed by a neural reconstruction process to build the initial 3D structure. This process is iteratively repeated to expand scene coverage by generating additional camera views and inpainting missing regions. However, such methods often suffer from prolonged generation times and incomplete coverage. Other studies have explored interactive and layout-guided approaches for 3D scene generation. Systems like WonderWorld (Yu et al., 2024a) provide user-friendly interfaces for scene customization, while Director3D (Li et al., 2024a) facilitates real-world scene creation with adaptive camera trajectories. Another potential limitation of this family of methods, as discussed in Sec. 1, is the lack of guaranteed global visual coherence, often resulting in distorted scene layouts and object shapes. Most recent literature has discovered the potential of video generation for consistent scene-level generation. VideoScene (Wang et al., 2025) leverages the temporal consistency in video diffusion models to generate novel views for 3D reconstruction by interpolating two input keyframes. Yu et al. and Wu et al. explore explicitly reconstructing a 3D point cloud as spatial memory for consistent video generation.
@@
The closest method to ours is DreamScene360 (Zhou et al., 2024), which generates 3D scenes from text prompts using a single panorama image, along with depth estimation and Gaussian Splats optimization. However, while the resulting scene supports 360° viewing, it suffers from a limited movable range. LayerPano3D (Yang et al., 2024) offers slightly increased navigational range by generating layered panoramas, but the content in the additional layers may suffer from distortion. ODIN (Wallingford et al., 2024) proposed learning 3D from 360° videos. However, their approach extracts only cropped views from videos to perform image-to-image style novel view synthesis, underutilizing the full potential of 360° videos. As a result, ODIN fails to model a complete 3D world and can only generate partial scenes. Beyond 3D, 4K4DGen (Li et al., 2024b), introduces panorama-to-4D generation but still remains limited in terms of traversable area.
In contrast to these approaches, our model introduces 360° videos as an explicit 3D scene proxy in the generation task, offering greater data scalability and ensuring strong visual coherence alongside comprehensive view coverage, ultimately enabling a realistic navigation experience.
3. WorldPrompter
3.1. Overview
@@
We show that 360° video capture serves as an effective, all-you-need proxy for 3D scene generation, as validated by the WorldPrompter pipeline. It synthesizes a traversable, 360° viewable 3D scene in two stages: generation and reconstruction. The generation stage leverages a text-conditioned 360° panoramic video diffusion model to generate a 10.6-second panoramic capture video for a target scene, simulating someone holding the camera and walking through the environment. Later, the reconstruction stage runs COLMAP (Schönberger and Frahm, 2016) to estimate precise camera poses from perspective crops of the generated videos and builds a navigable 3DGS using Long-LRM (Ziwen et al., 2024) with the calibrated camera poses. The generated panoramic video features rich details and geometrically consistent visual details over the entire scene, ensuring high-quality reconstruction and yielding appealing novel views in the traversable scene.
3.2. Video Generator for 360° Panoramic Captures